OpenStudy (anonymous):

I need a formula for population at t yrs when the population starts at 22mill and doubles in size every 7 yrs.

OpenStudy (anonymous):

\[22\times 2^{\frac{t}{7}}\] should work

OpenStudy (anonymous):

so t would be the percent it grows by each yr?

OpenStudy (anonymous):

no t is time in years

OpenStudy (anonymous):

fromula is \[P(t)=22\times 2^{\frac{t}{7}}\] that will give the population after "t" years so for example if t = 7,it should double as it is supposed to because \[P(7)=22\times 2^{\frac{7}{7}}=22\times 2=44\] as needed

OpenStudy (anonymous):

ok....I kept throwing the 22 and 2 around and just couldnt get them in the right places.

OpenStudy (anonymous):

thank you!